Fermi National Laboratory
 

Accelerator Update for December 11, 2006 - December 13, 2006

 

Monday December 11

The day shift began with Operations monitoring store 5116 (27.91E30) and a Recycler stash (170.7E10), with the Antiproton Source (Pbar) stacking (28.5E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 10:05 AM, Operations reported that the Pbar target had stopped rotating. Operations halted stacking while Pbar experts investigated.

At 10:49 AM, the MI-8 line suffered a vacuum burst, tripping off many vacuum pumps. Everything reset.

The Recycler completed a stash of antiprotons at 11:21 AM.

Operations prepared Pbar for access at 11:27 AM.

Operations terminated store 5116 at noon.

Cryo system technicians began wet engine repairs at TeV sectors A3 and C2 at 12:10 PM.

Operations prepared the CDF and D0 collision halls for an access at 12:19 PM. CDF came out of access at 1:52 PM. D0 came out a few minutes later.

Operations put the TeV through a dry squeeze at 2:54 PM.

Operations put the TeV into shot setup at 3:28 PM.

Pbar experts reported at 3:42 PM that they had completed their work on the target.

The Pbar access ended at 4:12 PM.

At 5:07 PM, Operations reported that stacking hadn’t resumed due to a problem with a power supply (A:IB). Operations called in an expert who repaired the supply.

Operations established store 5118 at 6:17 PM with an initial luminosity of 169.75E30. (MI did not suffer from NuMI beam fall out during TeV injections.)

The Recycler completed a stash around 9:50 PM.

Tuesday December 12

The midnight shift began with Operations monitoring store 5118 (74.7E30) and a stash (47.7E10), with Pbar stacking (48.8E10), and with NuMI and MiniBooNE taking beam.

Operations reported at 12:09 AM that operators had completed interlocking Meson enclosures 3, 4, and 5.

The Recycler completed a stash at 2:32 AM.

At 2:41 AM, two vacuum pumps (IP104 and 5) tripped off in the MI-10 area. The pumps reset okay.

At 2:45 AM, four MI-10 vacuum pumps tripped off (IP1041, 42, 51, & 52). They all reset except for 1041. 1042, 51, & 52 turned on, but 1051 tripped off immediately causing 1052 to trip. Eventually, operators left them all off. They will notify vacuum techs later in the morning of the problems.

At 2:58 AM, Operations reported that the H- Source had begun sparking about every five minutes.

The Recycler completed a stash at 4:26 AM.

At 5:34 AM, the Recycler electron cooling system tripped off. Operations tried various resets with no success and then called in experts.

At 6:19 AM, Operations had a problem with an Accumulator production readback; it was fluctuating between zero and large numbers. Operators tried various reboots without success. Operations called in experts.

The Recycler completed a stash at 7:24 AM.

Operations established beam to the Switchyard (SY) absorber at 8:39 AM.

At 9:06 AM, two MI kickers (I:KPS4A and B) tripped off on a “trigger ready” indication. A MI expert investigated. Beam resumed about thirty minutes later.

At 9:17 AM, a SY quadrupole power supply (S:Q202) wasn’t putting out any current. A power supply expert investigated and had the problem fixed 1.5 hours later.

At 10:10 AM, I:KPS4A and B tripped off again. A MI expert investigated and beam resumed about ten minutes later.

The Recycler completed a stash at 11:09 AM.

The Neutron Therapy Facility (NTF) reported at 1:11 PM that they would be conducting a beam study this afternoon.

Operations terminated store 5118 at 2:05 PM and then put the TeV into shot setup.

Operations reported at 3:27 PM that Booster beam was halted for RF experts to reconnect BRF2 into the system.

Operations established store 5119 at 4:18 PM with an initial luminosity of 159.85E30.

At 5:02 PM, BRF2 tripped off for the first of many, many trips in a little over an hour. An expert came in and made some adjustments after the last trip.

At 5:12 PM, LRF3 tripped off for the first of three times in ten minutes. It reset each time.

At 8:09 PM, a special MiniBooNE LCW system has been having trouble all afternoon and evening. The pump would turn on and seemingly come up to speed, but then trip off. Operations contacted an expert.

The Recycler completed a stash at 9:05 PM.

Operations reported at 11:43 PM that water experts haven’t been able to get the special MiniBooNE LCW system running. They will resume work on the system in the morning.

Wednesday December 13

The midnight shift began with Operations monitoring store 5119 (59.52E30) and the stash (148.9E10), with Pbar stacking (25.9E10), with MiniBooNE off due to LCW trouble, and with NuMI taking beam.

At 2:02 AM, Operations began tuning for a Pbar/Recycler stash, but ran into trouble. Operations decided to halt the stash and start over. The Recycler completed a stash at 3:23 AM.

Operations prepared the Meson M02 – 5 enclosures for access at 2:44 AM.

At 4:37 AM, the MI power supply loop tripped off on a “filter choke over current.” Operations reset the loop.

The Recycler completed a stash at 5:21 AM.

The Plans for Wednesday

The plans for today are to terminate store 5119 around 9 AM. MI experts will work on the MI abort kickers. The Run Coordinator will allow studies and maintenance work in all of the machines as long as the study or work won’t affect the Recycler stash.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/13/2006    email Fermilab